# SEMF: Supervised Expectation-Maximization Framework for Predicting Intervals

### v3

@@ Appendix E Full conformalized results
Interval Predictions Point Predictions
Relative Absolute Relative Absolute
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP MPIW NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
space_ga -8% 2% -11% 0.95 2.59 0.28 -13% -15% 0.55
space_ga -3% 0% -3% 0.95 2.77 0.30 -14% -17% 0.55
space_ga -2% 1% -4% 0.96 2.61 0.28 -13% -13% 0.57
space_ga -9% 1% -11% 0.96 2.81 0.30 -15% -18% 0.54
space_ga -7% 0% -8% 0.96 2.79 0.30 -19% -24% 0.51
cpu_activity 11% -4% 13% 0.95 0.57 0.11 -13% -20% 0.98
cpu_activity 14% -5% 16% 0.94 0.54 0.10 -13% -18% 0.98
cpu_activity 9% -4% 11% 0.95 0.58 0.11 -13% -18% 0.98
cpu_activity 10% -4% 13% 0.94 0.57 0.11 -15% -22% 0.97
cpu_activity 10% -4% 13% 0.94 0.57 0.11 -16% -22% 0.97
naval_propulsion_plant 131% 1% 56% 0.96 0.96 0.28 -320% -418% 0.96
naval_propulsion_plant 144% 1% 59% 0.96 0.88 0.26 -286% -369% 0.97
naval_propulsion_plant 174% 0% 63% 0.96 0.81 0.24 -257% -348% 0.97
naval_propulsion_plant 127% 0% 56% 0.96 0.94 0.28 -324% -426% 0.96
naval_propulsion_plant 108% 1% 52% 0.96 1.03 0.30 -391% -469% 0.95
miami_housing -7% 0% -7% 0.95 1.22 0.15 -11% -20% 0.90
miami_housing -10% -1% -10% 0.95 1.24 0.15 -6% -17% 0.90
miami_housing -8% -1% -8% 0.95 1.25 0.15 -13% -23% 0.89
miami_housing -9% -1% -10% 0.95 1.22 0.15 -9% -19% 0.90
miami_housing -9% 0% -9% 0.94 1.24 0.15 -10% -22% 0.89
kin8nm -10% 1% -12% 0.95 2.75 0.54 -36% -40% 0.46
kin8nm -8% 0% -9% 0.94 2.66 0.52 -33% -36% 0.49
kin8nm -11% 3% -15% 0.95 2.76 0.54 -32% -35% 0.50
kin8nm -9% 2% -12% 0.94 2.67 0.52 -36% -40% 0.48
kin8nm -10% 1% -13% 0.95 2.73 0.53 -33% -36% 0.49
concrete_compressive_strength -18% -3% -18% 0.89 1.67 0.34 -78% -110% 0.74
concrete_compressive_strength -5% -3% -3% 0.89 1.51 0.30 -71% -97% 0.76
concrete_compressive_strength -6% -6% 0% 0.87 1.42 0.29 -63% -87% 0.77
concrete_compressive_strength -6% -3% -4% 0.90 1.47 0.30 -63% -91% 0.77
concrete_compressive_strength -6% 1% -7% 0.92 1.59 0.32 -61% -88% 0.77
cars -35% 5% -60% 0.95 1.01 0.17 6% 3% 0.95
cars -25% 1% -34% 0.91 0.83 0.14 8% 4% 0.95
cars -19% 0% -24% 0.92 0.80 0.14 4% -2% 0.95
cars -24% -3% -28% 0.90 0.86 0.15 2% 0% 0.95
cars -21% -3% -22% 0.91 0.91 0.15 -3% 1% 0.94
energy_efficiency 16% -8% 21% 0.90 0.18 0.05 0% 0% 1.00
energy_efficiency 6% -2% 6% 0.97 0.21 0.06 4% 1% 1.00
energy_efficiency 8% -4% 11% 0.95 0.21 0.06 5% 2% 1.00
energy_efficiency 3% -3% 5% 0.96 0.21 0.06 1% 0% 1.00
energy_efficiency 17% -4% 18% 0.95 0.19 0.06 5% 0% 1.00
california_housing -2% -2% 1% 0.95 2.38 0.57 -16% -22% 0.70
california_housing 3% -2% 5% 0.95 2.31 0.55 -13% -19% 0.72
california_housing 2% -2% 4% 0.95 2.29 0.55 -15% -20% 0.71
california_housing 0% -3% 3% 0.95 2.32 0.55 -16% -22% 0.71
california_housing 1% -3% 4% 0.95 2.32 0.55 -14% -20% 0.72
airfoil_self_noise -25% -5% -26% 0.94 2.17 0.46 -166% -184% 0.74
airfoil_self_noise 0% -1% 1% 0.98 1.78 0.38 -59% -80% 0.89
airfoil_self_noise -15% -1% -16% 0.98 1.98 0.42 -97% -106% 0.84
airfoil_self_noise -28% -5% -32% 0.93 2.20 0.47 -169% -213% 0.67
airfoil_self_noise -14% -3% -13% 0.96 1.93 0.41 -98% -125% 0.84
QSAR_fish_toxicity -16% -2% -17% 0.88 2.45 0.38 -6% -10% 0.53
QSAR_fish_toxicity -10% -2% -9% 0.89 2.43 0.38 -7% -12% 0.52
QSAR_fish_toxicity 3% -6% 9% 0.88 2.38 0.37 -8% -12% 0.50
QSAR_fish_toxicity -12% -4% -9% 0.88 2.41 0.38 -5% -10% 0.55
QSAR_fish_toxicity 8% -9% 16% 0.84 2.03 0.32 -6% -13% 0.53

### v4

@@ Appendix E Full conformalized results on benchmark datasets
Relative Metrics Absolute Metrics
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW $\Delta$CRPS $\Delta$Pinball CWR NMPIW PICP
space_ga 5% -3% 8% 0% -8% 7.12 0.13 0.92
space_ga -6% -2% -4% -9% -13% 3.01 0.32 0.95
space_ga 3% 0% 3% 2% 3% 3.60 0.27 0.97
space_ga 25% -2% 21% 13% 8% 4.46 0.21 0.93
space_ga 0% 0% -1% 3% 8% 7.29 0.13 0.94
cpu_activity -16% -1% -18% -14% -10% 8.51 0.11 0.95
cpu_activity -9% 0% -11% -9% -11% 8.44 0.11 0.96
cpu_activity -14% 0% -16% -11% -11% 8.43 0.11 0.96
cpu_activity -16% 1% -20% -1% 17% 8.73 0.11 0.95
cpu_activity -14% 0% -16% -8% 0% 8.75 0.11 0.95
naval_propulsion_plant 84% -1% 46% 46% 33% 15.95 0.06 0.94
naval_propulsion_plant 26% 0% 20% 27% 0% 12.02 0.08 0.95
naval_propulsion_plant -2% -1% -1% -3% -25% 13.52 0.07 0.94
naval_propulsion_plant 8% 1% 7% 14% -25% 11.94 0.08 0.96
naval_propulsion_plant 1% 0% 1% 2% -20% 10.46 0.09 0.95
miami_housing -19% -1% -23% -13% -35% 9.00 0.11 0.95
miami_housing -35% 0% -56% -23% -38% 7.03 0.14 0.96
miami_housing -13% 0% -15% -9% -33% 9.98 0.09 0.94
miami_housing -20% 0% -25% -9% -11% 9.56 0.10 0.94
miami_housing -17% 0% -21% -9% -12% 9.71 0.10 0.95
kin8nm 17% 0% 14% 14% 15% 4.50 0.21 0.96
kin8nm 4% 1% 3% 5% 6% 4.86 0.20 0.95
kin8nm 18% -2% 16% 11% 11% 4.97 0.19 0.93
kin8nm 2% 0% 2% 5% 6% 4.59 0.21 0.95
kin8nm 3% 0% 3% 5% 0% 4.80 0.20 0.93
concrete_compressive_strength 54% -4% 37% 26% 21% 3.33 0.29 0.96
concrete_compressive_strength -5% -2% -2% -2% -19% 3.15 0.30 0.94
concrete_compressive_strength 13% -3% 14% 8% -12% 3.51 0.26 0.92
concrete_compressive_strength 73% -2% 43% 35% 31% 3.47 0.28 0.97
concrete_compressive_strength 19% 2% 14% 17% 15% 3.67 0.26 0.95
cars -11% -2% -11% -13% -29% 5.13 0.18 0.94
cars -7% -1% -7% -9% -13% 5.21 0.19 0.97
cars -18% 0% -22% -17% -19% 4.00 0.24 0.98
cars 21% -5% 22% 5% -21% 7.75 0.12 0.92
cars -44% 7% -89% -23% -14% 4.89 0.20 0.99
energy_efficiency 102% 5% 48% 56% 50% 19.46 0.05 0.96
energy_efficiency 74% -1% 42% 40% 40% 19.62 0.05 0.95
energy_efficiency 65% -3% 41% 36% 25% 18.91 0.05 0.96
energy_efficiency 46% 3% 30% 33% 0% 26.91 0.03 0.90
energy_efficiency 149% -6% 63% 42% 0% 24.65 0.04 0.92
california_housing -10% 0% -11% 0% -3% 1.99 0.48 0.96
california_housing -14% 0% -16% -5% -13% 2.12 0.45 0.94
california_housing -11% 0% -12% -1% -9% 2.12 0.45 0.95
california_housing -7% -1% -6% 0% -9% 2.22 0.42 0.94
california_housing -12% 0% -14% -2% -9% 2.20 0.43 0.95
airfoil_self_noise 60% -1% 38% 33% 30% 5.45 0.18 0.97
airfoil_self_noise 55% -1% 36% 37% 27% 4.83 0.20 0.97
airfoil_self_noise 63% 0% 39% 43% 32% 4.98 0.20 0.98
airfoil_self_noise 17% 1% 14% 20% 23% 4.84 0.20 0.95
airfoil_self_noise 16% -2% 15% 17% 5% 4.48 0.21 0.95
QSAR_fish_toxicity 21% -3% 19% 4% 6% 1.85 0.52 0.96
QSAR_fish_toxicity 2% 1% 1% 1% 2% 1.57 0.62 0.98
QSAR_fish_toxicity -2% 0% -2% -3% 2% 1.67 0.58 0.97
QSAR_fish_toxicity 36% -3% 28% 13% 15% 2.04 0.47 0.96
QSAR_fish_toxicity -15% 1% -19% -9% -11% 1.43 0.69 0.98